a sample of Freon -12 (CF2Cl2) occupies 10.0 L at 303 K and 166.70 kPa. Find it's volume at STP.

__________L
Answer #1

T1 = 303 K

P1 = 166.70 kPa

V1 = 10 L

At STP, : P2 = 101.325 kPa

T2 = 273.15 K

V2 = V (L)

According to ideal gas equation :

P1V1/T1 = P2V2/T2

166.7 kPa x 10 l /303 K = 101.325 kPa x V /273.15

5.502 = 0.371 V

V = 5.502/0.371 = 14.83 L

OR Gas law can be modified into

P1V1T2 = P2V2T1

166.7 kPa x 10 l x 273.15 K = 101.325 kPa x V x 303 K

455341.05 L/30701.475 = V

V = 14.83 L

Volume at STP = 14.83 L